# AugGen: Augmenting Task-Based Learning in Professional Creative Software with LLM-Generated Scaffolded UIs

### v1

@@ Abstract.
Professional creative software often presents steep learning curves due to complex interfaces, lack of structured task-aware guidance, and unfamiliar domain terminology. To address these challenges and augment user learning experience, we introduce AugGen, a method for generating scaffolded user interfaces that simplify interface complexity and support task-based learning. With the user’s task, our method surfaces task-relevant tools to reduce distracting features, organizes the tools around task workflow stages to offer execution guidance, connects tools with domain concepts to foster learning engagement, and progressively discloses advanced features to manage learning progress. To evaluate the method, we used our LLM-assisted pipeline to generate two task-specific scaffolded UIs and deployed them in Blender, our professional 3D modeling testbed. We invited both beginner (N=32) and expert (N=8) users to evaluate our implemented interfaces. Results show that the scaffolded interfaces significantly reduced user-perceived task load, enhanced task performance via embedded guidance, and augmented concept learning during task execution.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06
Figure 1. With AugGen, users can generate scaffolded user interfaces (UIs) to augment software and domain concept learning while executing their tasks in professional creative software. Taking a task description as input, AugGen automatically analyzes user task to decompose the task into workflow stages and extract relevant tools and domain concepts. The functionality of selected tools are implemented through programming and presented to the user as UI controls on the scaffolded interface. The code to implement functionality and UI controls are generated by LLMs and executed in the software for user interaction. Users can interact with the scaffolded interface with task-relevant tools structured around the workflow stages and learn domain concepts through labeled tooltips during task execution.
@@ 1. Introduction
Professional creative software, such as Photoshop (photoshop), Illustrator (illustrator), and Blender (blender), offers significant skill augmentation and forms the backbone of many creative and design industries. However, this skill comes at a cost of increased learning complexity. Prior work has identified that a major complexity arises from the user interface (UI), where newcomers often encounter hundreds of icons, menus, and specialized tools (chilana2018supporting). Learners frequently struggle to find the appropriate functions or tools and understand domain-specific terminology (novick2009micro; kiani2019beyond; lee2010usability). Compounding these issues is a lack of built-in structured guidance, which pushes users toward external resources, such as video tutorials, forums, and increasingly, large language models (LLMs), to navigate the UI and meet their task and learning goals. The continual toggling between disparate help sources disrupts workflow and frequently fails to deliver contextually appropriate support, thereby hindering self-directed learning and exploration (rieman1996field; chilana2012lemonaid; chilana2018supporting; khurana2024and).
@@
Multiple approaches have attempted to mitigate these challenges. One common strategy has been to provide simplified software alternatives, such as dedicated beginner modes (mcgrenere2002evaluation), separate novice-oriented applications with limited functionality (tinkercad; illustratoripad), or AI-generated task-specific interfaces (vaithilingam2024dynavis). While these alternatives reduce initial complexity, they often embody a simplicity-power tradeoff (uiandvis1999hearst), where users may plateau in simplified environments (cockburn2014supporting), or face a context switch when transitioning to the professional suite, which negates prior learning (nicol2024psychology; li2022design).
@@
Another line of work has focused on improving in-context assistance. Efforts range from static tooltips (grossman2010toolclips) and context-sensitive help (kelleher2005stencils) to more sophisticated systems that embed external resources directly into the application context (zhong2021helpviz). However, these approaches struggle with accurately inferring user goals, potentially leading to intrusive or irrelevant help (brdnik2022intelligent), and users may still find it challenging to map the provided information onto the complex software interface (kosch2023survey).
@@
More recently, the advancement of LLMs has led to the emergence of in-application software copilots that can process natural language queries, monitor visual context, or automate tasks (firefly; krol2025copilotvision; khurana2025me). While powerful for automating well-defined operations, current copilots struggle with complex, creative, or exploratory tasks (khurana2024and; qiao2024benchmarking). Furthermore, prior work has found that excessive automation or guidance by these copilots can reduce engagement (kosmyna2025your), diminish agency (li2024user; sellen2024rise), reduce the user’s ability to learn underlying concepts (chen2023understanding; vorvoreanu2025fostering), and contradict the well-established maxim of learning by doing (carroll1984training; rieman1996field). Lastly, excessive automation and prescriptive guidance from copilots can lead users to adopt a passive, instruction-following mode, akin to the studied “GPS effect” (ruginski2019gps), instead of learning.
@@
Despite substantial prior work, existing systems still fall short of holistically reducing UI complexity, offering in-context structured guidance, and helping users interpret domain-specific terminology in ways that integrate learning into task execution. Motivated by instructional-scaffolding research where temporary support helps learners build competence and gradually work independently (belland2017instructional), we introduce AugGen, a method that generates scaffolded interfaces using the underlying native software’s UI elements and operations. As shown in Figure 1, we demonstrate our LLM-based approach to generatively surface task-relevant tools from the underlying software and present them in a simplified layout tailored to the user’s current task. The resulting scaffolded UI progressively expands in complexity as the user gains proficiency, links interface elements with domain concepts to support concept learning, and guides the user through task workflows while remaining grounded in the native software environment.
@@
We apply our method to build two scaffolded interfaces in Blender (blender), a professional open-source 3D creation environment used as our testbed. We evaluate the two interfaces, one for 3D modeling and one for animation, with beginner (N=32) and expert (N=8) users. Results show that compared to the default Blender interface, the scaffolded UI has significantly improved task performance for both groups, as well as augmented concept learning for beginners. Based on a discussion of these findings, we present insights for future research on using our method to create scaffolded interfaces for instructional use cases to support both educators and learners, and to allow personalized workflows and cross-platform synchronization for end users. In summary, we make the following contributions:
@@
  • We propose a scaffolded interface generation method to reduce UI complexity of professional creative software, streamline task guidance through UI configuration, and augment domain concept learning during task execution.
@@ 2. Related Work
Our work builds upon prior research to address learning related challenges in professional creative software. Figure 2 presents a visual summary.
@@
Figure 2. Challenges in professional creative software (C1–C3), prior research to tackle these challenges, and AugGen’s focus.
@@
Challenges in professional creative software. Professional creative software has long been recognized as having a steep learning curve. A primary challenge is the complexity of the interface which presents numerous icons, tool palettes, and menus (chilana2018supporting). Prior work has identified that beginners frequently have trouble finding tools and determining the operational sequence for their tasks (novick2009micro; lee2010usability; kiani2019beyond). Another challenge is that the software provides limited guidance on how to proceed with user tasks. To address these challenges, users often turn to learning resources which are mainly external, e.g., formal manuals (novick2006don), tutorial videos (lafreniere2013community), blogs, and dedicated forums, which disrupt task workflows and create a high barrier to self-taught exploration (rieman1996field; chilana2012lemonaid; chilana2018supporting; khurana2024and). Additionally, professional creative software is designed for domain experts and its interfaces are often saturated with complex domain terminology, which raises difficulty and steepens the learning curve for beginners (popovic2000expert). In summary, the challenges of professional creative software that obstruct the learning experience include C1: complex UIs with extensive features; C2: limited context-aware, built-in guidance; C3: unfamiliar domain concepts.
Figure 3. AI copilots in Blender that offer step-by-step task instructions (left) (blender_copilot_oai) or automatically execute a task (right) (blender_copilot). The copilot’s instructions overlook interface cues and operational context. By assuming familiarity with Blender’s modes, shortcuts, and domain concepts, the guidance offers an incomplete procedural scaffold for beginner users.
@@
Simplified alternatives to professional creative software. To tackle these challenges, one approach has been to build software with simplified functionality for beginners, which addresses C1. For example, Adobe Express (adobeexpress), Canva (canva), OpenShot (openshot), TinkerCAD (tinkercad), and Illustrator on iPad (illustratoripad) target newcomers with simpler UIs and basic operations. In addition, recent work has explored natural language interfaces (generativefill; qin2024instructvid2vid; poole2022dreamfusion; liu2024dancegen) or LLM-generated interfaces (vaithilingam2024dynavis; liu2025crowdgenui), with simple interaction modalities like text or on-demand UI controls. These alternatives embody the classic simplicity-power tradeoff (uiandvis1999hearst), where simplified interfaces are less intimidating for novices, but they often lack the breadth of functionality. To balance this tradeoff, prior research has explored the transition between multiple interface modes, such as multi-layer interfaces (shneiderman2002promoting; mcgrenere2002evaluation) and workflow documentation interfaces (grossman2010chronicle).
@@
However, the gap between simple and advanced interfaces may lead to users getting stuck in the beginner mode if the two modes are too different (cockburn2014supporting), or getting overwhelmed if advancing to complex modes is poorly managed (nicol2024psychology). Another issue is that context switching from novice to professional software requires users to relearn workflows because the interfaces differ. Such discontinuity can void some of the initial learning done in the simplified application (li2022design). Our work aims to address these limitations by embedding a scaffolded interface into professional creative software to reinforce connection with the native software and reduce discontinuity from platform switching. To manage UI complexity and help users move from basic to advanced, we offer user-selectable complexity levels that let them gradually take on more advanced tools as they learn the underlying concepts and work through their tasks.
@@
In-context assistance for professional creative software. To address C2, research efforts have explored in-context assistance in professional creative software. Unlike external resources, such assistance aims to provide support at the user’s point of need, thus reducing disruption. For instance, previous work has designed elements, such as tooltips or help buttons, that are linked to relevant sections of a manual (farkas1993role; kelleher2005stencils; grossman2010toolclips; fourney2014intertwine; photoshop). While integrated into the application, these elements are static and often lack user context sensitivity. Recognizing the limitations of static help, researchers have developed context-aware approaches. One thread of work explores intelligent user assistance, such as using Bayesian probability networks to model user goals and offer help without explicit user requests (li2005active; horvitz2013lumiere). Another line of work focuses on improving contextual assistance delivery by bringing external resources into the application context, such as automatically presenting relevant web pages, documentation sections, forums, or tutorial videos based on the user’s current activity (chilana2012lemonaid; fraser2019replay; matejka2011ambient; matejka2011ip; ekstrand2011searching; zhong2021helpviz; yang2022softvideo; yang2024aqua).
@@
Nonetheless, existing in-context assistance faces challenges in accurately inferring user goals and risks providing irrelevant or poorly timed assistance (brdnik2022intelligent). In addition, while contextual assistance delivery reduces the need for explicit searching, users can still struggle to identify the most relevant information (lim2012improving) and map the provided guidance (textual steps or visual cues) onto the complex interfaces (kosch2023survey). We tackle these challenges by building a task-aware interface that surfaces the relevant tools from the underlying software to avoid contextually irrelevant support. Rather than asking users to hypothesize instructions and navigate a complex UI, our scaffolded interface embeds the guidance by grouping tools into workflow stages, aiming for a clear task-tool alignment.
@@
In-application software copilots. The recent advancement of LLMs has reshaped the landscape of in-context assistance, leading to the emergence of in-application AI copilots. These copilots can process natural language queries, offer instructions, and automate software tasks, e.g., copilots in Blender as shown in Figure 3. These copilots can also use multimodal information, such as screen monitoring, to understand user context and offer guidance, e.g., Microsoft Copilot Vision (krol2025copilotvision; cunningham2025copilotvision) and Google AI Studio (google2025aistudio). More recently, the copilots have unlocked the potential for on-demand UI generation (e.g., Gemini 3 (gemimi3)), which could be used to generate dynamic UIs to assist tasks based on user needs.
@@
Although these copilots are helpful to guide or automate tasks, excessive guidance and task automation can diminish the user’s sense of ownership over the task (endsley2017here; heer2019agency; li2024user; sellen2024rise) and reduce cognitive engagement in learning (kosmyna2025your). These copilots surface foundational concepts only incidentally, hindering a user’s conceptual understanding (C3) and running counter to the principle of learning by doing (carroll1984training; rieman1996field; novick2009micro; kiani2019beyond). The copilots might also suggest tools or UI elements that do not exist in the software or are not directly available, making the learning experience even more challenging. Although on-demand UIs can be generated by copilots to assist task execution, prior research has found that constantly changing UIs can increase cognitive load for the user, especially during longer task sessions (vaithilingam2024dynavis). To address these challenges and support durable learning, we embed explicit instructional scaffolds into the UI, aiming to engage the user in tasks. We bind each task-related action to its underlying concept to support learning while executing tasks. Furthermore, to reduce the cognitive load caused by dynamic UIs that may hinder learning, we offer users ready-to-use scaffolded UIs tailored to their tasks.
3. AugGen

### v2

# TaskLens: Generating Task-Conditioned Scaffolded Interfaces for Learning Professional Creative Software

@@ Abstract.
Professional creative software has steep learning curves for novices due to complex interfaces, limited guidance, and unfamiliar terminology. To support educators and tool creators in addressing learner challenges, we introduce TaskLens, an LLM-based method that automatically generates task-conditioned scaffolded UIs from natural language task descriptions. Our method uses LLMs to identify workflow stages and domain concepts, select task-relevant tools, generate implementation code, and execute the code to produce scaffolded interfaces. The interfaces surface relevant tools, organize them by workflow stage, link them to domain concepts, and progressively disclose advanced features. We evaluate TaskLens by deploying two LLM-generated scaffolded interfaces in Blender, a professional 3D modeling software. A user study with beginners (n=32) showed that our scaffolded interfaces significantly reduced perceived task load, improved task performance through embedded workflow guidance, and increased domain concept learning in Blender during task execution. A second study with experts (n=8) showed improved task efficiency and potential to create personalized UIs for productivity and creativity.
professional creative software, user interface generation, scaffolded learning, task assistance
journalyear: 2026copyright: ccconference: Designing Interactive Systems Conference; June 13–17, 2026; Singapore, Singaporebooktitle: Designing Interactive Systems Conference (DIS ’26), June 13–17, 2026, Singapore, Singaporedoi: 10.1145/3800645.3813081isbn: 979-8-4007-2563-0/2026/06ccs: Human-centered computing Interactive systems and tools
Figure 1. With TaskLens, users can generate task-conditioned scaffolded user interfaces (UIs) to augment software and domain concept learning while executing their tasks in professional creative software. Taking a task description as input, TaskLens automatically analyzes user task to decompose the task into workflow stages and extract relevant tools and domain concepts. The functionality of selected tools are implemented through programming and presented to the user as UI controls on the scaffolded interface. The code to implement functionality and UI controls are generated by LLMs and executed in the software for user interaction. Users can interact with the scaffolded interface with task-relevant tools structured around the workflow stages and learn domain concepts through labeled tooltips during task execution.
This figure shows a flowchart for scaffolded UI generation. Starting from the left, the flowchart takes user task as input, performs task analysis and domain concept extraction, generates UI implementation code with LLMs, and executes the code in professional creative software for user interaction.
@@ 1. Introduction
Professional creative software, such as Photoshop (Photoshop, 2025), Illustrator (Illustrator, 2025), and Blender (Blender, 2025d), offers significant skill augmentation and forms the backbone of many creative and design industries. However, this skill comes at a cost of increased learning complexity. Prior work has identified that a major complexity arises from the user interface (UI), where newcomers often encounter hundreds of icons, menus, and specialized tools (Chilana et al., 2018). Learners frequently struggle to find the appropriate functions or tools and understand domain-specific terminology (Novick et al., 2009; Kiani et al., 2019; Lee et al., 2010). Compounding these issues is a lack of built-in structured guidance, which pushes users toward external resources, such as video tutorials, forums, and increasingly, large language models (LLMs), to navigate the UI and meet their task and learning goals. The continual toggling between disparate help sources disrupts workflow and frequently fails to deliver contextually appropriate support, thereby hindering self-directed learning and exploration (Rieman, 1996; Chilana et al., 2012, 2018; Khurana et al., 2024).
@@
Multiple approaches have attempted to mitigate these challenges. One common strategy has been to provide simplified software alternatives, such as dedicated beginner modes (McGrenere et al., 2002), separate novice-oriented applications with simplified functionality (Autodesk, 2025; Adobe, 2025c), or AI-generated task-specific interfaces (Vaithilingam et al., 2024). While these alternatives reduce initial complexity, they often embody a simplicity-power tradeoff (Hearst, 1999), where users may plateau in simplified environments (Cockburn et al., 2014), or face a context switch when transitioning to the professional suite, which negates prior learning (Nicol, 2024; Li et al., 2022).
@@
Another line of work has focused on improving in-context assistance. Efforts range from static tooltips (Grossman and Fitzmaurice, 2010) and context-sensitive help (Kelleher and Pausch, 2005) to more sophisticated systems that embed external resources directly into the application context (Zhong et al., 2021). However, these approaches struggle with accurately inferring user goals, potentially leading to intrusive or irrelevant help (Brdnik et al., 2022), and users may still find it challenging to map the provided information onto the complex software interface (Kosch et al., 2023).
@@
More recently, the advancement of LLMs has led to the emergence of in-application software copilots that can process natural language queries, monitor visual context, or automate tasks (Adobe, 2025a; Krol, 2025; Khurana et al., 2025). While powerful for automating well-defined operations, current copilots struggle with complex, creative, or exploratory tasks (Khurana et al., 2024; Qiao et al., 2024). Furthermore, prior work has found that excessive automation or guidance by these copilots can reduce engagement (Kosmyna et al., 2025), diminish agency (Li et al., 2024; Sellen and Horvitz, 2024), reduce the user’s ability to learn underlying concepts (Chen et al., 2023; Vorvoreanu et al., 2025), and contradict the well-established maxim of learning by doing (Carroll and Carrithers, 1984; Rieman, 1996). Lastly, excessive automation and prescriptive guidance from copilots can lead users to adopt a passive, instruction-following mode, akin to the studied “GPS effect” (Ruginski et al., 2019), instead of critical thinking and learning (Swanson et al., 2026).
@@
Despite substantial prior work, existing systems still fall short of holistically reducing UI complexity, offering in-context structured guidance, and helping users interpret domain-specific terminology in ways that integrate learning into task execution. Motivated by instructional-scaffolding research where temporary support helps learners build competence and gradually work independently (Belland, 2017), we introduce TaskLens, a method that generates task-conditioned scaffolded interfaces to support software educators and tool creators in addressing the learner challenges using the underlying native software’s UI elements and operations. As shown in Figure 1, we demonstrate our LLM-based approach to generatively surface task-relevant tools from the underlying software and present them in a simplified layout tailored to the user’s current task. The resulting scaffolded UI progressively expands in complexity as the user gains proficiency, links interface elements with domain concepts to support concept learning, and guides the user through task workflows while remaining grounded in the native software environment.
@@
We apply our method to build two task-conditioned scaffolded interfaces in Blender [^1] https://www.blender.org/, a professional open-source 3D creation environment used as our testbed. We evaluate the two interfaces, one for 3D modeling and one for animation, with beginner (N=32) and expert (N=8) users. From the study with beginners (the software learners), we found that our scaffolded interfaces significantly reduced user-perceived task load, improved task performance through embedded workflow guidance, and increased domain concept understanding and learning in Blender during task execution. From the study with experts (the advanced users who have developed skills and knowledge through learning), we observed improved task efficiency and uncovered the potential to create personalized UIs for productivity and creativity needs of expert users. Based on a discussion of these findings, we present insights for future research on using our method to create scaffolded interfaces for instructional use cases to support both educators and learners, and to allow personalized workflows and cross-platform synchronization for end users. In summary, this work makes the following contributions:
@@
  • We propose a task-conditioned scaffolded interface generation method to reduce UI complexity of professional creative software, streamline task guidance through UI configuration, and augment domain concept learning during task execution.
@@ 2. Related Work
Our work builds upon prior research to address learning-related challenges in professional creative software. Figure 2 presents a visual summary.
@@
Figure 2. Challenges in professional creative software (C1–C3), prior research to tackle these challenges, and our focus.
This figure shows a venn diagram. There are three circles, each representing one of the three challenges this work addresses: C1 is complex UIs with extensive features, C2 is limited context-aware, built-in guidance, and C3 is unfamiliar domain concepts.
@@
Challenges in professional creative software. Professional creative software has long been recognized as having a steep learning curve. A primary challenge is the complexity of the interface, which presents numerous icons, tool palettes, and menus (Chilana et al., 2018). Prior work has identified that beginners frequently have trouble finding tools and determining the operational sequence for their tasks (Novick et al., 2009; Lee et al., 2010; Kiani et al., 2019). Another challenge is that the software provides limited guidance on how to proceed with user tasks. To address these challenges, users often turn to learning resources which are mainly external, e.g., formal manuals (Novick and Ward, 2006), tutorial videos (Lafreniere et al., 2013), blogs, and dedicated forums, which disrupt task workflows and create a high barrier to self-taught exploration (Rieman, 1996; Chilana et al., 2012, 2018; Khurana et al., 2024). Additionally, professional creative software is designed for domain experts and its interfaces are often saturated with complex domain terminology, which raises difficulty and steepens the learning curve for beginners (Popovic, 2000). In summary, the challenges of professional creative software that obstruct the learning experience include C1: complex UIs with extensive features; C2: limited context-aware, built-in guidance; C3: unfamiliar domain concepts.
@@
Simplified alternatives to professional creative software. To tackle these challenges, one approach has been to build software with simplified functionality for beginners, which addresses C1. For example, Adobe Express [^2] https://www.adobe.com/express/, Canva [^3] https://www.canva.com/, OpenShot [^4] https://www.openshot.org/, TinkerCAD [^5] https://www.tinkercad.com/, and Illustrator on iPad [^6] https://www.adobe.com/products/illustrator/ipad.html target newcomers with simpler UIs and basic operations. In addition, recent work has explored natural language interfaces (Adobe, 2025b; Qin et al., 2024; Poole et al., 2022; Liu and Sra, 2024) or LLM-generated interfaces (Vaithilingam et al., 2024; Liu et al., 2025), with simple interaction modalities like text or on-demand UI controls. These alternatives embody the classic simplicity-power tradeoff (Hearst, 1999), where simplified interfaces are less intimidating for novices, but they often lack the breadth of functionality. To balance this tradeoff, prior research has explored the transition between multiple interface modes, such as multi-layer interfaces (Shneiderman, 2002; McGrenere et al., 2002) and workflow documentation interfaces (Grossman et al., 2010).
@@
Nonetheless, the gap between simple and advanced interfaces may lead to users getting stuck in the beginner mode if the two modes are too different (Cockburn et al., 2014), or getting overwhelmed if advancing to complex modes is poorly managed (Nicol, 2024). Another issue is that context switching from novice to professional software requires users to relearn workflows because the interfaces differ. Such discontinuity can void some of the initial learning done in the simplified application (Li et al., 2022). Our work aims to address these limitations by embedding a task-conditioned scaffolded interface into professional creative software to reinforce connection with the native software and reduce discontinuity from platform switching. To manage UI complexity and help users move from basic to advanced, we offer user-selectable complexity levels that let them gradually take on more advanced tools as they learn the underlying concepts and work through their tasks.
@@
In-context assistance for professional creative software. To address C2, research efforts have explored in-context assistance in professional creative software. Unlike external resources, such assistance aims to provide support at the user’s point of need, thus reducing disruption. For instance, previous work has designed elements, such as tooltips or help buttons, that are linked to relevant sections of a manual (Farkas, 1993; Kelleher and Pausch, 2005; Grossman and Fitzmaurice, 2010; Fourney et al., 2014; Photoshop, 2025). While integrated into the application, these elements are static and often lack user context sensitivity. Recognizing the limitations of static help, researchers have developed context-aware approaches. One thread of work explores intelligent user assistance, such as using Bayesian probability networks to model user goals and offer help without explicit user requests (Li and Ji, 2005; Horvitz et al., 2013). Another line of work focuses on improving contextual assistance delivery by bringing external resources into the application context, such as automatically presenting relevant web pages, documentation sections, forums, or tutorial videos based on the user’s current activity (Chilana et al., 2012; Fraser et al., 2019; Matejka et al., 2011a, b; Ekstrand et al., 2011; Zhong et al., 2021; Yang et al., 2022, 2024), or showing procedural documentation with suggested next steps based on multiple prior demonstrations (Bergman et al., 2005).
@@
However, existing in-context assistance faces challenges in accurately inferring user goals and risks providing irrelevant or poorly timed assistance (Brdnik et al., 2022). In addition, while contextual assistance delivery reduces the need for explicit searching, users can still struggle to identify the most relevant information (Lim, 2012) and map the provided guidance (textual steps or visual cues) onto the complex interfaces (Kosch et al., 2023). Graphstract (Huang and Twidale, 2007) addresses the guidance-interface correlation challenge by using screen snippets centered around interaction areas to offer a visual overview of tools needed for user tasks. Although such graphic guidance can allow users to stay focused on task-related UI interactions, the task completion workflow remains implicit unless a detailed textual instruction is provided. In addition, for tasks that require using tools that are nested in menus or hidden in keyboard shortcuts and right-clicks, graphic guidance can become cluttered and ineffective. We tackle these challenges by building a task-aware scaffolded interface that surfaces the relevant tools, triggered by UI, keyboard, and mouse interactions, from the underlying software. Rather than asking users to hypothesize instructions and navigate a complex UI, our interface embeds the guidance by grouping tools into workflow stages to build a clear task-tool alignment. Crucially, most software assistance focuses on helping users complete tasks. Learning fundamental concepts during task performance has not been explicitly supported. We bridge this gap by integrating concept explanations into task instructions to actively engage users in learning concepts during tasks.
Figure 3. Examples of AI copilots that offer step-by-step task instructions (OpenAI, 2025) (left) or automatically execute a task (Blender, 2025c) (right) in Blender. The copilots’ instructions overlook interface cues and operational context. By assuming familiarity with Blender’s modes, shortcuts, and domain concepts, the guidance offers an incomplete procedural scaffold for beginner users.
This figure shows two examples of AI copilots in Blender. In the left figure, the copilot provides step-by-step task instructions, and in the right figure, the copilot automatically executes a user task.
@@
In-application software copilots. The recent advancement of LLMs has reshaped the landscape of in-context assistance, leading to the emergence of in-application AI copilots. These copilots can process natural language queries, offer instructions, and automate software tasks, e.g., copilots in Blender as shown in Figure 3. These copilots can also use multimodal information, such as screen monitoring, to understand user context and offer guidance, e.g., Microsoft Copilot Vision (Krol, 2025; Cunningham, 2025) and Google AI Studio (Google, 2025). More recently, the copilots have unlocked the potential for on-demand UI generation, e.g., generative and malleable UIs (Leviathan et al., ; Cheng et al., 2024; Cao et al., 2025; Min et al., 2025), which could be used to generate dynamic UIs to assist tasks based on user context and needs.
@@
Although these copilots are helpful to guide or automate tasks, excessive guidance and task automation can diminish the user’s sense of ownership over the task (Endsley, 2017; Heer, 2019; Li et al., 2024; Sellen and Horvitz, 2024), decrease user confidence during the task completion process (Xu et al., 2025), and reduce cognitive engagement in learning (Kosmyna et al., 2025). These copilots surface foundational concepts only incidentally, hindering a user’s conceptual understanding (C3) and running counter to the principle of learning by doing (Carroll and Carrithers, 1984; Rieman, 1996; Novick et al., 2009; Kiani et al., 2019). The copilots might also suggest tools or UI elements that do not exist in the software or are not directly available, making the learning experience even more challenging. Although on-demand UIs can be generated by copilots to assist task execution, prior research has found that constantly changing UIs can increase cognitive load for the user, especially during longer task sessions (Vaithilingam et al., 2024). To address these challenges and support durable learning, we embed explicit instructional scaffolds into the UI, aiming to engage the user in tasks. We bind each task-related action to its underlying concept to support learning while executing tasks. Furthermore, to reduce the cognitive load caused by dynamic UIs which may hinder learning, we offer users ready-to-use scaffolded UIs tailored to their tasks.
3. TaskLens